Find the total pressure ‘A’ at the bottom of the jar, 50 cm in height. The mass of the liquid is 4300 kg and the volume is 0.055 m3. The atmospheric pressure on Earth is 101.3 kPa.      






1
Expert's answer
2021-06-11T11:28:04-0400

density of liquid"=\frac{mass}{volume}=\frac{4300kg}{0.055m^3}=78181.82kg/m^3"

pressure due to liquid"=density\times gravity\times height=78181.82\times 9.8\times 50/100=383090.918Pa"

total pressure at the bottom"=383090.918+101,300=484,390.918Pa"